a line that passes through the point (-1,3) has a slope of 2. find onother point on the line.
Answer by sdmmadam@yahoo.com(530)


a line that passes through the point (-1,3) has a slope of 2. find onother point on the line.

Given slope and a point we use the slope and point formula to give the equation.
The slope and point formula is given by
(y-y1)/(x-x1) = m ----(*) where m is the slope and P(x1,y1) is the given point
Here P(x1,y1)=P(-1,3) and m =2
Therefore the required line is (using (*))
[y-(3)]/[x-(-1)]= 2
(y-3)/(x+1)=2
Multiplying by (x+1) through out,
y-3=2(x+1)
y-3 = 2x +2
0=2x-y+2+3 (taking all the terms from the left to the right, change side,then change sign)
0=2x-y+5
Answer: The required equation is 2x-y+5 = 0
Verification: putx = -1 and y = 3 in the above.
LHS = 2x-y+5 =2X(-1)-3+5 = -2-3+5 =0=RHS
Therefore our equation is correct.
Note:How is 0=2x-y+5 represented as
2x-y+5 = 0
Why sign has not been changed?
Answer:The equality relation is symmetric. That is p = q implies and is implied by q = p
Only one or two or three (...)terms transaction from one side to another side then change sign
Tom is as old as James is nothing but James is as old as Tom!